What is the pH of a 0.22-M solutio of carbonic acid? (Ka = 4.2 x 10^-7)

Please explain, thank you.
Answer #1

[H+] = sqrt(Ka*C)

       = sqrt(4.2*10^-7*0.22)

       = 0.000304

pH = -log(H+)

     = -log(0.000304)

     = 3.52